OpenStudy (anonymous):

The number of ordered pairs (P,Q) for which 213P43Q6 is divisible by 99 is

ganeshie8 (ganeshie8):

remember divisibility rule for \(9\) ?

OpenStudy (anonymous):

sum of all the digits should be a multiple of 9

ganeshie8 (ganeshie8):

yup! setup a equation

ganeshie8 (ganeshie8):

\(2+1+3+P+4+3+Q+6 = 9k \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

but it is 99

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

if 99 divides a number, then 9 also has to divide that same number right ?

ganeshie8 (ganeshie8):

so, we're not doing anything illegal.. go ahead setup the equation !

OpenStudy (anonymous):

19 + P + Q =9k

ganeshie8 (ganeshie8):

yes, we need another equation

ganeshie8 (ganeshie8):

since 99 = 9*11, lets use divisibility rule for 11

ganeshie8 (ganeshie8):

setup another equation

OpenStudy (anonymous):

so 11k right?

OpenStudy (anonymous):

19 + P + Q =11k ?

ganeshie8 (ganeshie8):

\(2-1+3-P+4-3+Q-6 = 11k \)

ganeshie8 (ganeshie8):

http://www.mathsisfun.com/divisibility-rules.html

ganeshie8 (ganeshie8):

you need to take alternating sum for 11

OpenStudy (anonymous):

oh so i will simplify wait

OpenStudy (anonymous):

-1 - P + Q = 11k

ganeshie8 (ganeshie8):

yes, so we have two equations : 1 + P + Q =9k -1-P+Q = 11m

OpenStudy (anonymous):

yes but how to solve them simultaneously we have three variables

ganeshie8 (ganeshie8):

easy, we dont solve the varibales

ganeshie8 (ganeshie8):

just see that P and Q are just "digits" that means they have just SINGLE CHARACTER right ? they have to be between 0-9

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

1 + P + Q =9k -1-P+Q = 11m

ganeshie8 (ganeshie8):

look at first equation, we want "1 + P + Q" to be multiple of 9

ganeshie8 (ganeshie8):

that means, P+Q can be 8 or 17

ganeshie8 (ganeshie8):

youw wid me ?

OpenStudy (anonymous):

how 17

ganeshie8 (ganeshie8):

1 + P + Q =9k plugin k = 1, 2, ...

ganeshie8 (ganeshie8):

we dont want values from k=3 because they make each of P or Q greater than 9.

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

P + Q = 17 P + Q = 8

OpenStudy (anonymous):

^ these are the equations

ganeshie8 (ganeshie8):

# 1 + P + Q =9k you got : P + Q = 17 P + Q = 8 #-1-P+Q = 11m you got : ??

ganeshie8 (ganeshie8):

find the equations for 11m also

OpenStudy (anonymous):

How many numbers to put in because it is -P

ganeshie8 (ganeshie8):

good question, max value of any digit = 9 min value of any digit = 0 so the difference should not exceed 9

ganeshie8 (ganeshie8):

-1-P+Q = 11m put m = 0

OpenStudy (anonymous):

Q-P=1

ganeshie8 (ganeshie8):

So, it all boils down to solving : # 1 + P + Q =9k you got : P + Q = 17 P + Q = 8 #-1-P+Q = 11m you got : -P + Q = 1

ganeshie8 (ganeshie8):

take the first equation from first group take the first equation from second group solve them

ganeshie8 (ganeshie8):

P+Q = 17 -P + Q = 1

ganeshie8 (ganeshie8):

wat do u get when u solve them ?

OpenStudy (anonymous):

(8,9)

ganeshie8 (ganeshie8):

bingo ! thats ur one solution ! lets see if we can get one more solution or not

OpenStudy (anonymous):

we won't

ganeshie8 (ganeshie8):

take the second equation from first group take the first equation from second group and solve them

ganeshie8 (ganeshie8):

how do u say that ?

ganeshie8 (ganeshie8):

P + Q = 8 -P + Q = 1

ganeshie8 (ganeshie8):

solve them

OpenStudy (anonymous):

Because the answer is only (P,Q)

OpenStudy (anonymous):

I mean it is given in the book

ganeshie8 (ganeshie8):

solve them and see wat u get ?

OpenStudy (anonymous):

But logically it seems correct to solve the equations but we don't get whole numbers

ganeshie8 (ganeshie8):

exactly ! you got it !!

OpenStudy (anonymous):

we don't get whole numbers

OpenStudy (anonymous):

u genius man!! india se hai kya

ganeshie8 (ganeshie8):

for below equations : P + Q = 8 -P + Q = 1 we wont get whole numbers for P and Q, so this is not a valid solution. throw it !

ganeshie8 (ganeshie8):

only one ordered pair exists (P, Q) = (8, 9)

OpenStudy (anonymous):

Good problem thank u you

ganeshie8 (ganeshie8):

lol yes india... but no genius lol

ganeshie8 (ganeshie8):

you're the genius :)

OpenStudy (anonymous):

kidhar mumbai me rehta hai

ganeshie8 (ganeshie8):

bangalore, wbu ?

OpenStudy (anonymous):

mumbai

OpenStudy (anonymous):

what is ur name?

OpenStudy (anonymous):

ganesh tadi i got it

ganeshie8 (ganeshie8):

ganesh yes :) btw wat u studying ?

ganeshie8 (ganeshie8):

ahh ur profile is empty, wats ur name ?

OpenStudy (anonymous):

I am preparing for IIT . I passed tenth last year in november it is because i am from different board not sse. So i have joined resonance this year and my entrance test is on 30 th march .

OpenStudy (anonymous):

Udit kulkarni

OpenStudy (anonymous):

So u an IIT

OpenStudy (anonymous):

i mean iitian

ganeshie8 (ganeshie8):

omg! i thought you were atleast 12th + ! lol you're too smart !!

ganeshie8 (ganeshie8):

nope, and im done wid my studies, been working from > 5 years actually...

OpenStudy (anonymous):

So what is IC engineeing like never heard of it actually before

ganeshie8 (ganeshie8):

gtg... need to get ready for work... good luck wid entrance exam :) im sure u wil do great !! cya :)

ganeshie8 (ganeshie8):

chip designing

ganeshie8 (ganeshie8):

you must have heard about "integrated circuits" ?

OpenStudy (anonymous):

Thank you and wish u luck too!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool!!

ganeshie8 (ganeshie8):

IC design involves designing those chips wil talk more ltr byeee have a great day !

OpenStudy (anonymous):

same 2 u bye